Two friends rent an apartment while they go to college for $975 per month. Since one bedroom is 60 square feet larger than the other bedroom, each person's rent contribution is based on bedroom size. Each person agrees to pay $3.25 per square foot of bedroom area. Let x be the area (in square feet) of the larger bedroom, and let y be the area (in square feet) of the smaller bedroom. Write and solve a system of linear equations to find the area of each bedroom. I need a step by step not just an answer because I don't get how to start the equations off.
Answer by mananth(16946)

x-y =60..............1
3.25(x+y)=975
/3.25
x+y = 300........2
add (1) & (2)
2x=360
/2
x=180 sq. ft. Larger bedroom
120 sq.ft smaller bedroom